# A Survey on Causal Discovery: Theory and Practice

### v1

@@ 5.5 Interventional Algorithms
$\mathcal{S}_{\mathcal{I}}(G)=\sup_{\phi}\sum_{\textbf{I}\in\mathcal{I}}\mathbb{E}_{X}\log f^{(\textbf{I})}(X;B,R,\phi)-\lambda|G|$ (5.14)
where $\phi$ are the weights of the estimator used to maximize the score function (i.e. neural networks in this case), $X$ follows the interventional distribution $P^{(\textbf{I})}$, $f^{(\textbf{I})}$ is the interventional density function, $B$ the binary adjacency matrix of $G$, $R$ the binary interventional matrix (i.e. $R_{ij}=1$ if $X_{i}\in\textbf{I}_{j}$).
Essentially, the score function is built upon the conditional interventional distribution to recover the invariant edges across interventions. In fact, vertices that are not in any intervention target are characterized by a conditional probability distribution which is invariant across interventions, as for conservative families of interventions. Relying on conditional invariance, the causal graph $\hat{G}=\operatorname*{\arg\!\max}_{G\in\mathbb{G}}\mathcal{S}_{\mathcal{I}}(G)$ is $\mathcal{I}$-equivalent (Subsection 5.8) to the true graph $G^{*}$, for $\lambda>0$ small enough. In case of unknown intervention targets, an additional $-\lambda_{R}|\mathcal{I}|$ regularization term is added to the score function.
The DCDI algorithm has been tested against IGSP and GIES with known interventions and JCI-PC and UT-IGSP for unknown interventions, showing marginal advantages in terms of structural recovery. As for others continuous optimization methods [114], the major strength point is represented by its scalability: it takes $O(n^{3})$ to compute the matrix exponential during each training step, making it the only causal discovery algorithm that supports non-linear relationships in the interventional setting in a high-dimensional setting.
@@ 6 Evaluation and Tuning
This section tackles the evaluation and tuning step typical of any practical application. A collection of reference data sets is listed, both real-world and synthetic generated ones, serving as benchmarking resources for discovery methods. In order to evaluate different solutions resulting from a set of configurations (i.e. hyperparameter) we report comparison metrics found in the specialized literature, both in terms of structure and entailed causal statements. Finally, tuning strategies and software packages are explored as support for new developed techniques.
@@ 6.1 Evaluation Datasets
Ever-growing data set [62, 105] designed to benchmark discovery algorithms against bi-variate settings with known ground truth. The latest version reported by the change-log (December 20, 2017) includes 108 pairs. Each pair is composed by a data file with two columns (cause and effect respectively), a short description of the data sampling procedure, and a 2D scatter plot.
@@
Flow cytometry measurements [84] of 11 proteins and phospholipids. The data set is split in different experiments, with nine stimulatory or inhibitory interventions. The study compares new learned model against ground truth obtained by reference literature on signaling networks with intervention points.
@@
Single-cell RNA-sequencing (scRNA-seq) data set [44] of $\sim$3000 mouse embryonic stem cells after leukemia inhibitory factor (LIF) withdrawal. The ground truth model is obtained by querying the TRRUST database [30] for the related causal relationships.
@@
The DREAM4 challenge [92] provides five datasets simulated from five biologically plausible gene regulatory networks with 10 genes [54]. Each dataset is composed by both observational and interventional data sampled by applying small random noise perturbations, single-gene knockdowns and single-gene knockouts, resulting in time series with unknown interventions.
Dataset Type URL
Tuebingen [62] Real-world Here
CausalWorld [2] Synthetic Here
Sachs [84] Real-world Here
Klein [44] Real-world Here
Perturb-Seq [21] Real-world Here
SynTReN [110] Synthetic Here
DREAM4 [92] Synthetic Here
Table 4: Static datasets by type and availability.
@@ 6.2 Evaluation Metrics
In the context of causal discovery, the definitions of true positive (TP), true negative (TN), false positive (FP) and false negative (FN) have the same interpretation of the metrics referred to a binary classifier, which try to predict the edge orientation.
@@
A first set of evaluation metrics for graphical models is made of the adjacency precision (AP) and adjacency recall (AR) [85]. These metrics are computed as the ratio between the number of correctly predicted adjacent vertices over the total predicted ones for AP and true predicted ones for AR. Formally, once the confusion matrix associated with the presence of edges is computed, the two metrics are defined as follows:
$AP=\frac{TP}{TP+FP}$ (6.1) $AR=\frac{TP}{TP+FN}$ (6.2)
@@
6.3 Parameter Tuning
Strategies to perform parameter/s turning are rarely found in surveys, even tough casual discovery algorithms may have multiple parameters that regulate the search procedure. Here, we report three general and flexible practices described in the specialized literature that can be applied to any technique described so far.
@@ 6.3 Parameter Tuning
A first approach for parameter/s tuning is related to model complexity. The goal is to find the parameter/s configuration that minimizes the complexity of the associated causal graph. As a measure of complexity one can rely on the Bayesian Information Criterion (BIC) [87] or the Akaike Information Criterion (AIC) [3]. This tuning strategy is particularly effective when coupled with score-based approaches that are able to exploit the same function, allowing to reuse the intermediate scores for a faster evaluation. The most general form of model complexity minimization is implemented as a grid search over all parameters configurations for the ranges.
@@
Out-of-sample Causal Tuning (OCT & OCTs)
While previous approaches focused on metrics related to the causal structure alone, authors in [10] propose to employ the resulting model for its prediction capabilities, reducing the problem into an evaluation of a predictor. This approach works in a out-of-sample fashion, hence the name Out-of-sample Causal Tuning (OCT). The main advantages of such method are (i) the lack of parametric assumptions about the distribution of the data and (ii) the generalization to cases where the BIC and AIC scores are not defined, i.e. discrete models with hidden variables.
@@ 6.4 Software Packages
The Causal Discovery Toolbox [42] is a Python front-end that acts as a bridge between different subpackages, pooling together multiple discovery algorithms. For example, one may find constrained-based algorithms such as PC, Max-Min Parents & Children (MMPC) [109], score-based algorithms as GES and variants (GIES), and non linear approaches as LiNGAM, Causal Additive Models (CAM) [14], and others.
@@
While previous packages where intended for command line usage, TETRAD [76] is a causal discovery package developed in Java with a graphical user interface. It follows a pipeline paradigm, where the user can drag & drop boxes from the side bar and connect them together to form data pipelines. The user can choose from a wide range of options, such as PC (PCStable, CPC & CPCStable [18], PCMax), FCI (RFCI, RFCI-BSC [39], GFCI), GES (FGES, FGES-MB, IMaGES), LiNGAM, and others. Given the simplicity of interface, it is well suited for researcher with limited programming experience.
@@ 7.2 Causal Discovery in Medicine
The methods were executed both with and without trivial background knowledge, e.g. patient’s age is not affected by any biomarker. A significant improvement was observed with the addition of the knowledge base. Finally, longitudinal data were included, discovering more edges and removing the incorrect ones. The performance of the constraint-based was lower and less stable across the bootstrap samples than the score-based one.
@@ 7.3 Causal Discovery in Psychology
Psychopathology researchers in [7] used graphical modeling algorithms to identify causal relationships within and between manifestations of psychiatric disorders. In this context, such methods are employed to identify symptoms that are part of a causal chain of “mediators”. The main target of the study was to test whether drinking motivated by the goal of reducing negative affect (i.e. drinking to cope, DTC) served as a mediator in comorbid alcohol use and anxiety disorder in a causally insufficient setting.
@@ 8.2 Future Directions
In terms of opportunities for future extensions, in this contribution we did not explore the implications of applying such method to time series, which would add an additional layer of complexity. Indeed, the representation of the causal dependencies in time are different from the one expressed in a static scenario and deserves a separated discussion on its own, especially when combined with the other topics introduced during the discussion.

### v2

@@ 5.5.7 Differentiable Causal Discovery with Interventions (DCDI)
$\mathcal{S}_{\mathcal{I}}(G)=\sup_{\phi}\sum_{\textbf{I}\in\mathcal{I}}\mathbb{E}_{X}\log f^{(\textbf{I})}(X;\textbf{B},\textbf{R},\phi)-\lambda|G|$ (5.14)
where $\phi$ are the weights of the estimator used to maximize the score function (i.e. neural networks in this case), $X$ follows the interventional distribution $P^{(\textbf{I})}$, $f^{(\textbf{I})}$ is the interventional density function, B the binary adjacency matrix of $G$, R the binary interventional matrix (i.e. $R_{ij}=1$ if $X_{i}\in\textbf{I}_{j}$) and $\lambda$ a penalty coefficient.
Essentially, the score function is built upon the conditional interventional distribution to recover the invariant edges across interventions. In fact, vertices that are not in any intervention target are characterized by a conditional probability distribution that is invariant across interventions, as for conservative families of interventions. Relying on conditional invariance, the causal graph $\hat{G}=\operatorname*{\arg\!\max}_{G\in\mathbb{G}}\mathcal{S}_{\mathcal{I}}(G)$ is $\mathcal{I}$-equivalent (Subsection 5.8) to the true graph $G^{*}$, for $\lambda>0$ small enough. In case of unknown intervention targets, an additional $\lambda_{R}|\mathcal{I}|$ regularization term is added to the score function.
The DCDI algorithm has been tested against IGSP and GIES with known interventions and JCI-PC and UT-IGSP for unknown interventions, showing marginal advantages in terms of structural recovery. As for others continuous optimization methods [47], the major strength is represented by its scalability: it takes $O(n^{3})$, with $n$ the number of variables, to compute the matrix exponential during each training step, making it the only causal discovery algorithm that supports non-linear relationships in the interventional setting in a high-dimensional setting.
@@ 6 Evaluation and Tuning
This section tackles the evaluation and tuning step typical of any practical application. A collection of reference datasets is listed in Table 4, both real-world and synthetic generated ones, serving as benchmarking resources for discovery methods. In order to evaluate different solutions resulting from a set of configurations (i.e. hyperparameters) we report comparison metrics found in the specialized literature, both in terms of structure and entailed causal statements. Finally, tuning strategies and software packages are explored as support for new developed techniques.
@@ 6.1 Evaluation Datasets
Ever-growing dataset [95, 45] designed to benchmark discovery algorithms against bi-variate settings with known ground truth. The latest version reported by the change-log (December 20, 2017) includes 108 pairs. Each pair is composed by a data file with two columns (cause and effect, respectively), a short description of the data sampling procedure, and a 2D scatter plot.
@@
Flow cytometry measurements [97] of 11 proteins and phospholipids. The dataset is split into different experiments, with nine stimulatory or inhibitory interventions. The study compares new learned model against ground truth obtained by reference literature on signaling networks with intervention points.
@@
Single-cell RNA-sequencing (scRNA-seq) dataset [98] of mouse embryonic stem cells after leukemia inhibitory factor (LIF) withdrawal. The ground truth model is obtained by querying the TRRUST database [99] for the related causal relationships.
@@
The DREAM4 challenge [102] provides five datasets simulated from five biologically plausible gene regulatory networks with 10 genes [103]. Each dataset is composed by both observational and interventional data sampled by applying small random noise perturbations, single-gene knockdowns and single-gene knockouts, resulting in time-series with unknown interventions.
Dataset Source Type URL
Tuebingen [95] R O Here
CausalWorld [96] S I Here
Sachs [97] R M Here
Klein [98] R M Here
Perturb-Seq [100] R I Here
SynTReN [101] S O Here
DREAM4 [102] S M Here
Table 4: Static datasets by source, type and availability. Each dataset originates from (R)eal-world or (S)ynthetic experiments, which may contain (O)bservational, (I)nterventional or (M)ixed data, i.e. both observational and interventional instances.
@@ 6.2 Evaluation Metrics
In the context of causal discovery, the definitions of true positive (TP), true negative (TN), false positive (FP) and false negative (FN) have the same interpretation of the metrics referred to a binary classifier, which tries to predict the edge orientation.
@@
A first set of evaluation metrics for graphical models is made of the adjacency precision (AP) and adjacency recall (AR) [104]. These metrics are computed as the ratio between the number of correctly predicted adjacent vertices over the total predicted ones for AP, and true predicted ones for AR. Formally, once the confusion matrix associated with the presence of edges is computed, the two metrics are defined as follows:
$AP=\frac{TP}{TP+FP}$ (6.1)
$AR=\frac{TP}{TP+FN}$ (6.2)
@@
6.3 Parameters Tuning
Strategies to perform parameters tuning are rarely found in surveys, even tough casual discovery algorithms may have multiple parameters that regulate the search procedure. Here, we report three general and flexible practices described in the specialized literature that can be applied to any technique described so far.
@@ 6.3 Parameters Tuning
A first approach for parameters tuning is related to model complexity. The goal is to find the parameters configuration that minimizes the complexity of the associated causal graph. As a measure of complexity one can rely on the Akaike Information Criterion (AIC) (Equation 3.5) or the Bayesian Information Criterion (BIC) (Equation 3.6). This tuning strategy is particularly effective when coupled with score-based approaches that are able to exploit the same function, allowing to reuse the intermediate scores for a faster evaluation. The most general form of model complexity minimization is implemented as a grid search over all parameters configurations for the ranges.
@@
Out-of-sample Causal Tuning (OCT)
While previous approaches focus on metrics related to the causal structure alone, authors in [107] propose to employ the resulting model for its prediction capabilities, reducing the problem into an evaluation of a predictor. This approach works in a out-of-sample fashion, hence the name Out-of-sample Causal Tuning (OCT). The main advantages of such method are i) the lack of parametric assumptions about the distribution of the data, and ii) the generalization to cases where the BIC and AIC scores are not defined, i.e. discrete models with hidden variables.
@@ 6.4 Software Packages
The Causal Discovery Toolbox [108] is a Python front-end that acts as a bridge between different subpackages, pooling together multiple discovery algorithms. For example, one may find constraint-based algorithms such as PC, Max-Min Parents & Children (MMPC) [77], score-based algorithms as GES and variants (GIES), and non linear approaches as LiNGAM, Causal Additive Models (CAM) [109], and others.
@@
While previous packages were intended for command line usage, TETRAD [116] is a causal discovery package developed in Java with a graphical user interface. It follows a pipeline paradigm, where the user can drag & drop boxes from the side bar and connect them together to form data pipelines. The user can choose from a wide range of options, such as PC (PCStable, CPC & CPCStable [39], PCMax), FCI (RFCI, RFCI-BSC [117], GFCI), GES (FGES, FGES-MB, IMaGES), LiNGAM, and others. Given the simplicity of the interface, it is well suited for researchers with limited programming experience.
@@ 7.1 Causal Discovery in Economics
Using PC for multi-variate analysis in an economic application lead to a high-level understanding of the influences present in the global economic cycle. Moreover, authors highlight the advantages of applying causal discovery approaches w.r.t. other methodologies, e.g. “Granger causality”, especially with the selected algorithms. In fact, the PC algorithm is specifically designed to avoid conditioning on irrelevant variables, resulting in larger effect size and lower dimensionality.
@@ 7.2 Causal Discovery in Medicine
Moreover, performing the discovery step alternatively under causal sufficient and insufficient assumptions provides useful hints to isolate the influence of unmeasured external variables, e.g. the presence of a bi-directed edge instead of a directed one. These differences in the recovered structures guide researches during the interpretation of the collected data, and ease the evaluation of competing hypotheses that are consistent with the causal graph.
The methods were executed both with and without trivial background knowledge, e.g. patient’s age is not affected by any biomarker. A significant improvement was observed with the addition of the knowledge base. Finally, longitudinal data were included, discovering more edges and removing the incorrect ones. The performance of the constraint-based approach was lower and less stable across the bootstrap samples than the score-based one.
@@
Researchers explicitly selected this approach in order to rule-out direct and indirect effects of latent variables, while taking advantage of the hybrid structure of the algorithm to speed-up the discovery step. In fact, the GFCI algorithm leverages an initial phase that exploits the score-based approach described in FGES, which inherently caches intermediate evaluations of the score function to reduce the execution time.
@@ 7.3 Causal Discovery in Psychology
Psychopathology researchers in [121] used graphical modeling algorithms to identify causal relationships within and between manifestations of psychiatric disorders. In this context, the GFCI algorithm was employed to identify symptoms that are part of a causal chain of “mediators”. The main target of the study was to test whether drinking motivated by the goal of reducing negative affect (i.e. drinking to cope, DTC) served as a mediator in comorbid alcohol use and anxiety disorder.
One of the reasons why GFCI is well suited for this scenario is due to the capability of this algorithm to tackle the causally insufficient assumption. This property constitutes a major advantage, enabling researches to explore the effect of potential hidden variables during the assessment of the drinking-anxiety interplay in general.
@@ 8.2 Future Directions
In terms of opportunities for future extensions, in this contribution we did not explore the implications of applying such method to time-series, which would add complexity. Indeed, the representation of the causal dependencies in time are different from the one expressed in a static scenario and deserves a separated discussion on its own, especially when combined with the other topics introduced during the discussion.
From a more general perspective, there is a set of issues that have yet to be solved in order to shorten the distance between these techniques and real-world applications. For instance, causal discovery in presence of missing data is an under explored field. In fact, imputing missing data when the missingness mechanism is not at random results in sub-optimal solutions due to the introduction of bias component. There have been some recent advances w.r.t. to this topic [122, 123], but these new contributions are restricted to constraint-based approaches only.
Another aspect that is worth mentioning is the application of causal discovery to heterogeneous data [124]. The “heterogeneity” of the data is referred to the collection process that happens under different observational contexts. Naïvely pooling different data sources together into a single one results in poor performances. Essentially, this problem stems from the presence of distribution shifts, which arise when context variables differ significantly, e.g. in case of environmental changes. Still, learning a causal representation in the presence of multiple heterogeneous sources is a relevant issue that has yet to be solved.
Finally, there are other scenarios in which causal discovery is limited by the current state-of-the-art, such as non-identical overlapping sets of variables [28, 125], learning from streaming data [126] and federated learning [127].
Acknowledgements
We thank the two anonymous reviewers for their insightful comments. We also thank professor Daniela Besozzi for her thoughtful suggestions.